Workkng alone Carrie can paint a house in 3 hours less than Walter. Together Carrie and Walter an paint a house in 8 hours. How long does it take Walter to paint a house?
Answer by edjones(8007)

Walter can paint the house in w hours; Carrie in w-3 hrs.
1/w the amt. walt does in 1 hr.; 1/(w-3) carrie does in 1 hr.
1/w + 1/(w-3) = 1/8 together in 1 hr.
8(w-3)+8w=w(w-3)
8w-24+8w=w^2-3w
w^2-19w+24=0 Quadratic formula
w=17.6 hrs
.
Ed